2. Starting from the formula for gravitational potential energy: Uo -S LMA, derive the universal law of gravitation: Pe 3. Starting from the definition of the resultant force being the time derivative of translational momentum, show that the resultant force is only equal to ma if 0. Hint: Use the product rule.
Answer #1

a)

gradient of the potential anergy gives the force,

F=-\frac{dU_G}{dx}

Putting potential energy in the above term will get,

F=-\frac{d}{dr}\left ( \frac{Gm_1m_2}{r} \right )

Finding the differentiation,

F=\frac{Gm_1m_2}{r^2}

b)

Here force is given by,

F=\frac{dp}{dt}

where p is the momentum,

F=\frac{d(mv)}{dt}

Using the product rule,

F=m\frac{d(v)}{dt}+v\frac{d(m)}{dt}

Form above it is very much clear that F=ma is the valid expression if and only if we have dm/dt =0.

F=ma \left [ \because \frac{dm}{dt}=0 \right ]